OpenStudy (anonymous):

Find the slope of the line. 4x+5y=3

OpenStudy (anonymous):

solve for y first 5y = 3 - 4x y= (3-4x)/5 y = 3/5 - 4/5x the number next to the x is your slope so your slope is -4/5